OpenStudy (anonymous):

graph the question y=8x-7

OpenStudy (anonymous):

the y=mx+b form give the y-intercept (0,-7) to get the x-int we can let y=o and solve for x: 0=8x-7 7=8x x=7/8 so the other point is (7/8, 0)

OpenStudy (anonymous):

:})

OpenStudy (anonymous):

thans